OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 4 to 9. The perimeter of the larger rectangle is 54 inches. What is the perimeter of the smaller rectangle? A. 36 inches B. 13.5 inches C. 24 inches D. 121.5 inches

OpenStudy (moneybaby980):

A

OpenStudy (trojanpoem):

small/ big = 4/9 small/54 = 4/9 can you continue ?